

********************************************************************
UARS WEEKLY STATUS REPORT - 28 August 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM has successfully gathered its normal solar, offset, and occulation
data during this past week without apparent problems.
***********
PEM
***********
PEM was not active
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
HALOE Status
August 28, 1998
HALOE was turned on at the beginning of day today. All engineering
parameters are nominal. Science channel performance is being evaluated,
but initial look indicates they are working properly.
***********
MLS
***********
The instrument is now off as scheduled.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
95 days HAL (4.80) PROD (L1) (1873R-1232C-25F=616L)
44 days HAL (4.80) PROD (L3) (1873R-714C=1159L)
80 days SOL (4.62) PROD (L2-L3) (2349R-1247C=1102L)
2 days SUS (7.00) ENG (2R-2C=0L)
The CDHF is preparing for the upgrade of current FORTRAN compiler,
version 6.3, to version 6.5. A statement of impact has been
has been sent to the user community. Plans to install, following
response to user correspondence, were scheduled for the
week of the 24th of August, however network problems took precedence.
The CDHF worked with NSI personnel and the UARS Network Manager
toward the resolution of network problems affecting entire gateway.
CDHF pending software upgrades: IDL 5.1, OpenIngres, Fortran V6.5.
***********
FOT
***********
REPORTING PERIOD
This report covers 15 August 1998 (Orbit 37846, GMT Day 227) through 21
August 1998 (Orbit 37950, GMT Day 233).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 227/0112 37860 AXIS Off
HRDI 230/2240 37905 Low Power
WINDII 231/0024 37906 On
HALOE 233/1524 37945 Off
The beta angle ranged between the angles of 46.4 degrees to 16.5 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 0.0 degrees on 25 August 1998 (DOY 237).
Five (5) additional SSPP SOLSTICE Offset sequence were performed this week
on GMT days 228-232 at the request of SUSIM operations (and with the
agreement of SOLSTICE operations). The offsets were identical to the
regular SSPP SOLSTICE Offset sequence, which is nominally scheduled once
per week.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 14 August 1998 and 21 August 1998 are provided in
Appendix D.
The clock error ranged from 10.2 to -12.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
- ALERTS = 01
- OUT-OF-LIMITS = 05
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
- FOT 03
- HRDI 02
- HALOE 01
- PEM 01
- WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
- FOT 04
One (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/210 - 2287.496941
Transponder B: 98/211 - 2287.499161
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/232 show the observatory data loss to be
21 hours, 04 minutes, 45.332 seconds (an increase of eighteen (18) minutes
and 34.784 seconds since last report period). This is a 0.0348 percent
data loss which equals a 99.9652 percent data capture for the mission.
The increase of 18 minutes and 34.784 seconds occurred over six days:
- 18:29.472 seconds on Day 227 due to insufficient overlap and degraded
data;
- 1.216 seconds on Day 228 due to insufficient overlap and line errors;
- 2.592 seconds on Day 229 due to line errors and hits in data;
- 0.160 seconds on Day 230 due to line errors;
- 1.120 seconds on Day 231 due to line errors; and
- 0.224 seconds on Day 232 due to line errors.
Applications Processor (AP) ground system software testing has been
suspended waiting release of an operational pack containing version 10.11.
FOT performed an operational pack "cleanup" this week in order to
facilitate a baseline the new Ground Software using the current operational
pack. Redundant OBC dumps on the pack were eliminated keeping the required
one set per month. This resulted in a substantial disk space saving.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
The FOT is investigating an incident and maybe a second incident of high
noise at the beginning of playback of the Narrowband Tape Recorder (NBTR)
in an attempt to identify whether it is due to flight equipment or ground
equipment.
Flight operations is investigating an alternate to sun pointing safe hold
which can provide better protection for a two battery power subsystem. The
FOT is also attempting to define a one battery scenario which would save
the spacecraft and allow a single instrument to operate or the SSPP
instruments to operate with all other instruments completely OFF.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) No change in status.
User Planning System (UPS) - No change in status.
Multi Satellite Operations Control Center (MSOCC) - No change in status.
Command Management System (CMS) - The FOT worked with CMS Year 2K Test Team
representatives Daya Srivastava and Sidney Kirschner this week in an
continuing effort to get the back-up CMS system ready for testing. Due to
an increased disk quota for the test account on UCMSB, the FOT was able to
continue copying necessary support files from the prime CMS to the back-up
system. In addition, the FOT worked with Mr. Kirschner in attempting to
retrieve FDF Year 2K planning aids from a 4mm tape generated in the FORMATS
Test Facility. Due to incompatibilities between the system that was used
to generate the tape and the CMS system, the files could not be downloaded
onto UCMSB. The FOT and the Year 2K test team are investigating other
alternatives to getting the FDF data onto the back-up UCMS system.
There was no Year 2K CMS status meeting this week.
Flight Dynamics Facility (FDF) - Year 2K planning aids have been generated.
OBC PMON FLIGHT SOFTWARE UPDATE
Flight Software - The Flight Software is complete.
Ground Support Software - Complete. The Applications Processor operational
disk pack is being generated.
Verification Testing - Testing of the operations pack compilation of the
software will be required prior to uplink.
IP NETWORK TESTING
Space Network (SN) - NCC I/P testing was conducted during TDRSS contact at
232/15:22 to 225/15:36. Initially, no NCC data and no acknowledgments
GCMRUs were received. Later in the contact, good acknowledgment for our
GCMRUs were received but ODM data was present. The problem is currently
under investigation and additional testing will be scheduled.
Testing with the NCC 98 system to support routine flight operations is
being planned.
Deep Space Network (DSN) and Ground Spaceflight Tracking and Data Network
(GSTDN) - Working with the DSN test lead for I/P testing to verify UARS
emergency backup sites.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
The B/D thruster refill effort continues. Latch valve testing at Eaton has
been completed and preliminary results given to the GSFC Propulsion Branch.
A more comprehensive Eaton report is expected to follow. The ground line
fill tests are still planned to be executed after analysis of the Eaton
data.
Jackson and Tull were authorized to go forward on the electronics for latch
valve operation pending release of further funds.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing support to the re-engineering effort to
process the UARS telemetry on the TOMS/EP control center computer system.
UARS COMMAND RE-HOST PHASE 1 DEMONSTRATION
Flight operations continues to support the update of the Lockheed Martin
SCS-21 control center system to support a phase 1 feasibility
demonstration. FOT provided a cross reference of ODM block data to the ODM
display page.
The SCS-21 equipment arrived at GSFC on 17 August 1998 and is being
assembled and operationally checked.
NCC 98 availability for SCS-21 testing is limited to a single 8-hour-block
during the week of 24 August 1998.
MEETINGS
The FOT met with the Project Scientist and Deputies on 19 August 1998 to
brief them on the UARS subsystem status. Various contingencies and
possible reaction were also discussed.
No UARS Battery telecon was held this week. The next telecon has been
scheduled for 25 August 1998 at 1:00 pm.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 25 August
1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 August 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
37848 227/0431 98-298 OBC Clk Adj fr 20.6 to 0.00msec/day
37860 227/0112 98-299 PEM AXIS Power Off
37880 229/0637 98-300 OBC Clk Adj fr 0.00 to 20.6msec/day
37905 230/2240 98-301 OBC Clk Adj fr 20.6 to 0.00msec/day
37905 230/2240 98-302 HRDI Low Power Mode
37906 231/0024 98-303 WINDII Power ON
37944 233/1415 98-304 HRDI Htr. Cycle to 75 amp-min
37945 233/1524 98-305 HALOE Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 August 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-HI WINDII EU Temperature 37868,73,74,876-906
HRPTSPBIMON Y-HI PTS Pulse Bus Cur. Mon 37892
HRIFESURVTMP Y-HI IFE Survival Temp 37907
HRPSEQBIMON Y-HI HRDI PSE Q/B Cur. Mon. 37942-4
QTSISAIF Y-HI SSPP Baseplate SISA I/F 37942-4
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WISRCBBS Broad Band Source Op (during Turn-on) 37907
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 August 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
37848 227/04:38 (15 Aug 98) from 20.6 to 00.0 msec/day
37880 229/06:43 (17 Aug 98) from 00.0 to 20.6 msec/day
37905 230/22:42 (18 Aug 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 August 1998
14 Aug 1998 - GMT Day 226 Beta = 46.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.1 0.000 0.0
2 +22.4/-22.4 +6.11/+4.51 +28.4/-16.0 25.1 1.024 21.6
3 +33.6/-16.8 +5.15/+3.23 +27.2/-16.0 25.0 1.026 21.1
21 Aug 1998 - GMT Day 233 Beta = 16.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/ +0.0 18.1 0.000 0.0
2 +67.2/-0.0 +7.09/+4.83 +31.2/-16.4 23.8 1.023 23.1
3 +39.2/-33.6 +5.79/+3.87 +30.4/-16.8 23.8 1.023 22.4
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 August 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-047 FOT 37862 228/0253 14170 Generic Late Acquisition
98-048 FOT 37870 228/1555 N/A RCDR in PLBK at LOS
98-050 FOT 37932 232/1840 N/A degraded 512 kbps dump
98-051 FOT 37938 223/0459 N/A TAC N0-8 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-046 FOT 37742 220/0116 20129 TELEMETRY DROPOUT
98-048 FOT 37870 228/1155 NBR NOT IN REC
98-049 FOT 37831 226/0114 NOISY NBR P/B
98-050 FOT 37932 232/1840 NOISY NBR P/B
98-051 FOT 37938 233/0500 TAC 8 CRASH
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-045 FOT See list Aug 1998 14170 Generic Late Acquisition
98-047 FOT 37856 227/1713 20149 NOISY NBR P/B